OpenStudy (anonymous):

(6x^3 + 18x^2) - 3x^2 =

OpenStudy (anonymous):



OpenStudy (anonymous):

you want this simplified right?

OpenStudy (anonymous):

if so, all your doing is combining like forms... here are the steps of doing so... (6x^3 + 18x^2) - 3x^2 = Remove parentheses...6x^3 + 18x^2 - 3x^2 = now if you look you will see that 18x^2 and 3x^2 are like terms sooo you combine to get 15x^2 that leaves you with it simplified at 6x^3+15x^2 Hope this helps!

OpenStudy (anonymous):

=6x^3 + 18x^2 - 3x^2 =6x^3+15x^2 = 3x^2 (2x+5)

OpenStudy (anonymous):

I would just combine like terms. first the 18x^2 -3x^2 for which you get 15x^2. since you cant do anything with the 6x^3, because the exponets are different then you have 6x^3+15x^2 :D